OpenStudy (anonymous):

what number decreased by 1/3 of itself equals 32

OpenStudy (anonymous):

A

OpenStudy (bibby):

\(\large x-\frac{1}{3}x=32\)

OpenStudy (anonymous):

Let the number be : \(x\) It is decreased by 1/3 of itself: that means : \(\large x - \frac{1}{3}x\) Equals 32 means : \(= 32\) So, finally you have : \(\large x - \frac{1}{3}x = 32\) Try to solve for \(x\) here..